Please answer quickly!!! Recently, More Money 4U offered an annuity that pays 5.7% compounded monthly. If $1,189 is deposited into this annuity every month, how much is in the account after 9 years? How much of this is interest?

Type the amount in the account: $____

Type the amount of interest earned: $____

Please do not round any numbers.

(Please square, label, or circle the answer)
Answer #1

It is assumed payment is made at beginning of month
Number of payments = 9 x 12 = 108
Interest per period = 5.7%/12 = 0.475%

Amount in account = FV(5.7%/12,108,1189,0,1)

= $168072

Amount of Interest earned = $168072 - 1189 x 108 = $39660

For payment made at end of month

Amount in account = FV(5.7%/12,108,1189,0,0)

= $167278

Amount of Interest earned = $167278 - 1189 x 108 = $38866